OpenStudy (anonymous):

Evaluate S5 for 600 + 300 + 150 + … and select the correct answer below.

OpenStudy (welshfella):

this is another geometric series:- r will be second term / first term

OpenStudy (anonymous):

What would the equation for this one be?

OpenStudy (solomonzelman):

you can use the formula \(\large\color{black}{ \displaystyle {\rm Sum} _{\rm ~(when~~adding~~some~~n~~number~~of~~terms)} \displaystyle =\frac{({\rm 1st~ term)\times (1-{\rm r}^n)}}{1-{\rm r}}}\)

OpenStudy (welshfella):

r = 300 / 600 = ?

OpenStudy (anonymous):

37.5 is the answer